# On Touchard’s Identity: Generalizations and Related Results

### v2

@@ 2 A generalization of Touchard’s identity and related results
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{-2k}}}{{2k+r}}\,C_{k}}=\frac{{1}}{{n+1}}\binom{{n+r}}{{n+1}}^{-1}\sum_{k=0}^{n}{2^{-k}C_{k+1}\binom{{k+r-1}}{k}}$. (27)
@@
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{-2k}}}{{4k-1}}\,C_{k}}=2^{2n}\binom{{2n}}{n}^{-1}\left({2\sum_{k=1}^{n}{2^{-3k}C_{k-1}C_{k+1}}-1}\right).$ (30)
@@
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}H_{2k+r}}=2^{-n}C_{n+1}H_{n+r}-\sum_{k=0}^{n-1}{\binom{{n}}{k}\frac{2^{-k}}{n-k}C_{k+1}\binom{{n+r}}{{k+r}}^{-1}}.$ (50)
@@
and take the limit as $s$ approaches zero, using (4). ∎
@@ 4 Related results via binomial transform
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{2\left({n-2k}\right)}}}{{n-2k+1}}\,C_{k}H_{n-2k}}=\frac{1}{2}\sum_{k=0}^{\left\lceil{n/2}\right\rceil}{\binom{{n}}{{2k-1}}\frac{{2^{2k}}}{k}\,C_{n-2k+2}O_{k}}.$ (81)

### v3

# arXiv:2602.10558v3[math.CO] 01 Mar 2026

@@ 2 A generalization of Touchard’s identity and related results
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{-2k}}}{{2k+r}}\,C_{k}}=\frac{{1}}{{n+r}}\binom{{n+r-1}}{{n}}^{-1}\sum_{k=0}^{n}{2^{-k}C_{k+1}\binom{{k+r-1}}{k}}$. (27)
@@
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{-2k}}}{{4k-1}}\,C_{k}}$ $=2^{2n}\binom{{2n}}{n}^{-1}\left({2\sum_{k=1}^{n}{2^{-3k}C_{k-1}C_{k+1}}-1}\right),$ (30)
@@
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{-2k}}}{{2k-1}}\,C_{k}}=n\left({-1+\sum_{k=2}^{n}{2^{-k}\,\frac{{C_{k+1}}}{{k\left({k-1}\right)}}}}\right).$ (31)
@@
To prove (31), write (27) as
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{-2k}}}{{2k+r}}\,C_{k}}$ (35)
$\qquad=\frac{1}{{n+r}}\left({\left({1+r}\right)\binom{{n+r-1}}{n}^{-1}+\sum_{k=2}^{n}{2^{-k}C_{k+1}\binom{{k+r-1}}{k}\binom{{n+r-1}}{n}^{-1}}}\right),$ (36)
and use
$\lim_{r\to-1}\left({1+r}\right)\binom{{n+r-1}}{n}^{-1}=-n\left({n-1}\right)$ (37)
and
$\lim_{r\to-1}\binom{{k+r-1}}{k}\binom{{n+r-1}}{n}^{-1}=\frac{{n\left({n-1}\right)}}{{k\left({k-1}\right)}},$ (38)
since
$\left({1+r}\right)\binom{{n+r-1}}{n}^{-1}=\frac{{\left({n+r}\right)\left({n+r+1}\right)}}{r}\binom{{n+r+1}}{n}^{-1}$ (39)
and
$\binom{{k+r-1}}{k}\binom{{n+r-1}}{n}^{-1}=\frac{{n+r}}{{k+r}}\binom{{n}}{{n+r}}\binom{{k}}{{k+r}}^{-1}.$ (40)
@@
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}H_{2k+r}}=2^{-n}C_{n+1}H_{n+r}-\sum_{k=0}^{n-1}{\binom{{n}}{k}\frac{2^{-k}}{n-k}C_{k+1}\binom{{n+r}}{{k+r}}^{-1}}$ (56)
and
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{-2k}}}{{2k+r}}\,C_{k}H_{2k+r-1}}$ (57)
$\qquad=\frac{1}{{n+r}}\binom{{n+r-1}}{n}^{-1}\sum_{k=0}^{n}{2^{-k}C_{k+1}\left({H_{k+r-1}-H_{n-k}}\right)\binom{{k+r-1}}{k}}$. (58)
@@
and take the limit as $s$ approaches zero, using (4). Identity (58) is obtained by setting $s=1$ in (55) and making use of (32). ∎
@@
Proposition 8.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{-2k}}}{{2k+1}}\,C_{k}H_{2k}}=\frac{1}{{n+1}}\sum_{k=0}^{n}{2^{-k}C_{k+1}\left({H_{k}-H_{n-k}}\right)}$ (69)
and
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{-2k}}}{{4k+1}}\,C_{k}O_{2k}}=\frac{{2^{2n-1}}}{{2n+1}}\binom{{2n}}{n}^{-1}\sum_{k=0}^{n}{2^{-3k}C_{k+1}\binom{{2k}}{k}\left({2O_{k}-H_{n-k}}\right)}.$ (70)
Proof.
Identity (69) is the evaluation of (58) at $r=1$. Identity (70) is obtained by plugging $r=1/2$ in (58), making use of (16) and (33) and equating rational parts. ∎
@@ 4 Related results via binomial transform
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{2\left({n-2k}\right)}}}{{n-2k+1}}\,C_{k}H_{n-2k}}=\frac{1}{2}\sum_{k=1}^{\left\lceil{n/2}\right\rceil}{\binom{{n}}{{2k-1}}\frac{{2^{2k}}}{k}\,C_{n-2k+2}O_{k}}.$ (91)